Find the equilibrium point for the demand and supply functions for
D(p)=133-10p
S(p)=65+7p

Found 2 solutions by Fombitz, Edwin McCravy:
Answer by Fombitz(32388)
Answer by Edwin McCravy(20054)

The equilibrium point is the value of the price p when demand function D(p)
equals the supply function S(p):

So set D(p)=S(p) by setting their right sides equal:

133-10p = 65+7p

and solve for the price p (which makes demand equal to supply.

Edwin